Given he population of numbers 3,6,8,9 and 4. Suppose samples of size 3 are drawn from this population .


1. What is the mean and the variance of the population?

1
Expert's answer
2022-03-19T02:41:18-0400

"EX=\frac{3+6+8+9+4}{5}=6\\DX=\frac{\left( 3-6 \right) ^2+\left( 6-6 \right) ^2+\left( 8-6 \right) ^2+\left( 9-6 \right) ^2+\left( 4-6 \right) ^2}{5}=5.2"